### v1

# Fermi-LAT Discovery of a Gamma-ray Outburst from the Peculiar Compact Steep Spectrum Radiogalaxy 3C 216

Abstract
3C 216 is an extra-galactic radio source classified as a compact steep spectrum (CSS) object, associated with the source 4FGL J0910.0+4257 detected by the Large Area Telescope (LAT) on board the Fermi Gamma-ray Space Telescope. The source exhibits extended radio structures as well as an inner relativistic jet. In general, jets accelerated by active galactic nuclei (AGNs) are efficient sources of non-thermal radiation, spanning from the radio band to X-ray and gamma-ray energies. Due to relativistic beaming, much of this radiation, particularly in the high-energy domain, is concentrated within a narrow cone aligned with the jet’s direction. Consequently, high-energy emission is more easily detected in blazars, where the jet is closely aligned with the line of sight of the observer. Beginning in November 2022 , Fermi-LAT observed increased gamma-ray activity from 3C 216, culminating in a strong outburst in May 2023. This event was followed up by observations from the Neil Gehrels Swift Observatory telescope. In this work, we perform a careful analysis of the multifrequency data (gamma-ray, X-ray, UV, optical) collected during this observational campaign. We find that the spectral energy distribution of the flaring source evolves in a coherent way, suggesting a common origin for the multifrequency emission. These results support the interpretation of the gamma-ray emission within a single zone synchrotron self-Compton (SSC) model, with important implications for the mechanisms powering high-energy radiation in AGN jets.
gamma rays: galaxies — gamma-ray astronomy — high energy astrophysics — gamma-ray sources: individual (3C 216) — radio galaxies: individual (3C 216) — X-ray sources: individual (3C 216) — blazars — relativistic jets — spectral energy distribution
journal: The Astrophysical Journal
1 Introduction
The active galactic nucleus (AGN) 3C 216 ($z=0.670$, (Smith & Spinrad, 1980)) is classified as an extragalactic CSS radio source. On arcsecond scales, it consists of a central component surrounded by a more extended structure with an angular size of $4.5^{\prime\prime}$, corresponding to a projected linear size (LS) of $56$ kpc (Principe et al., 2021). Although its radio spectrum peaks at low frequency ($\nu<0.5$ GHz), an upturn is observed at a few GHz, attributed to the presence of a central flat-spectrum core (Taylor et al., 1995). This core exhibits a significant misalignment with the outer structure, along with superluminal jet component motion, with a velocity of approximately $\sim 4c$ (Venturi et al., 1993; Paragi et al., 2000). While the compact nature of CSS sources is typically attributed to their young age within an evolutionary framework (Fanti et al., 1995), the characteristics of the central component, a pronounced optical polarization and variability (Impey et al., 1991), strongly suggest the presence of a blazar core.
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope, is a pair-conversion telescope designed to detect photons in the energy range from $20$ MeV to $2$ TeV. The LAT observed a pronounced enhancement in the gamma-ray activity from the direction of the source on 2022-12-08 (La Mura, 2022) and on 2023-05-01 (Giacchino et al., 2023). Multifrequency observations of the source have been carried out in order to associate the gamma-ray flare with activity at different wavelengths. The observations across various bands provide insights into both the origin of the gamma-ray emission and the structure of the radio jet. As we will demonstrate in this work, during the flare emission, the source exhibited a hard gamma-ray spectrum, in contrast to its typically soft spectrum during quiescent periods. This behavior can be explained by a synchrotron self-Compton (SSC) model across all frequencies, with evidence of a cooling process in the days after the flare peak.
In the Sec. 2 we report the main results of the Fermi-LAT observations of the flaring activity that occurred in May 2023 and in the Subsec. 2.1 we provide a refined analysis of the Fermi-LAT data, highlighting the main features of the source in terms of its gamma-ray spectrum and light curve. We then describe in Sec. 3 the Swift ToO observations of 3C 216, with a detailed analysis presented in Subsec. 3.1. Finally, the discussion and conclusion are provided in Sec. 4.
2 Fermi-LAT observations
The CSS source 3C 216 is located at R.A. = 137.38957 deg, DEC. = 42.89624 deg [J2000,  (Petrov et al., 2005)]. It is associated with the 4FGL J0910.0+4257 source in the third data release of 4FGL catalog (4FGL-DR3, Abdollahi et al., 2022) with coordinates R.A. = $137.51\pm 0.11$ deg, DEC. = $42.96\pm 0.11$ deg.
Fermi-LAT detected flaring activity from the direction of this source. A preliminary Fermi-LAT analysis of the Flare Advocate data for 2023-05-01 revealed that the daily-averaged gamma-ray flux increased by a factor of $177$ relative to the average flux reported in the 4FGL-DR3, $\langle\Phi\rangle_{\gamma}=(7.5\pm 1.4)\times 10^{-9}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ in the $1-100$ GeV energy range. The photon index decreased from the 4FGL averaged value, ($2.57\pm 0.10$), to ($2.11\pm 0.09$). This represented both the highest daily flux and the hardest photon index recorded for this source by Fermi-LAT to date (Giacchino et al., 2023). Due to the exceptional magnitude of this flare, the source association has been confirmed through a dedicated follow-up analysis with Fermi-LAT, complemented by multifrequency observations with Swift, which we will describe in detail later.
2.1 Gamma-ray analysis
The data analysis was performed with Fermipy [^1] https://fermipy.readthedocs.io/en/latest/ (version 1.2.0.) (Wood et al., 2017) and the Fermitools [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software packages (version 2.2.0). We analyzed three periods: all Fermi-LAT data up to one month after the event (see Table 1 for details), the flare period, defined as the epoch during which Fermi-LAT detected an increasing activity trend on a weekly time scale, and the peak flare (PF) period, covering the maximum daily activity and the subsequent multi-frequency follow-up granted by the requested Swift observations.
Parameter name Value
Time domain (Gregorian) $2008/08/04$ to $2023/06/06$
Energy range 100 MeV to 300 GeV
IRF P8R3_SOURCE_V3
Event Type FRONT + BACK
Point Source Catalog 4FGL-DR3
ROI size $15^{\circ}\times 15^{\circ}$
Pixel size $0.1^{\circ}$
Bins per energy decade 8
Galactic diffuse model gll_iem_v07.fits
Isotropic diffuse model P8R3_SOURCE_V3_v1.txt
Table 1: Table of Fermipy analysis parameters
The time range chosen for flare period is $2022$-$11$-$14$ to $2023$-$06$-$06$, while the PF period goes from $2023$-$04$-$28$ to $2023$-$05$-$11$. The energy range selected for the analysis is $[0.1-300]$ GeV, using Pass 8 events (Atwood et al., 2013) and all the available photons of the SOURCE class, excluding those arriving with zenith angles greater than $90^{\circ}$ for energies smaller than $1$ GeV otherwise greater than $105^{\circ}$. Moreover, we choose a region of interest (ROI) of $15^{\circ}\times 15^{\circ}$ around our target, with a pixel size of $0.1^{\circ}$ and $8$ evenly spaced logarithmic energy bins. We utilize the P8R3_SOURCE_V3 instrumental response functions (IRFs), along with the galactic diffuse model gll_iem_v07.fits and the isotropic diffuse model P8R3_SOURCE_V3_v1.txt. The model used for the analysis includes all sources in the 4FGL-DR3 catalog [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html located at a distance $\leq 20^{\circ}$ from 4FGL J0910.0+4257. A summary of the parameters is reported in Table 1. In the spectral analysis we left the model parameters of the isotropic and diffuse background as well as the sources within $3^{\circ}$ of our target free to vary. To produce the spectral energy distribution (SED) of all the three periods, we used eight logarithmically spaced bins between $100$ MeV and $10$ GeV and four selected energy bins from $10$ GeV and $300$ GeV obtaining the plot in Fig. 1. The plot clearly shows that the PF (blue) and flare (green) period fluxes are much higher than the values obtained considering all the Fermi-LAT (red) observations, confirming an interesting enhanced activity.
Figure 1: SED of 3C 216 for three periods: in blue PF (Peak Flare) period, in green flare period, in red total Fermi-LAT period. The upper limit is reported when $TS\leq 10$
Period $\Gamma$ $\Phi_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ $\alpha$ $\beta$ $N_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$
PF $1.97\pm 0.05$ $(5.87\pm 0.46)\times 10^{-11}$ $1.79\pm 0.23$ $0.12\pm 0.09$ $(3.83\pm 0.72)\times 10^{-11}$
Flare $2.03\pm 0.01$ $(2.45\pm 0.09)\times 10^{-11}$ $1.99\pm 0.16$ $0.06\pm 0.03$ $(1.73\pm 0.25)\times 10^{-11}$
Total $2.17\pm 0.04$ $(1.69\pm 0.07)\times 10^{-12}$ $1.95\pm 0.20$ $0.12\pm 0.07$ $(1.02\pm 0.47)\times 10^{-12}$
DR4 $2.43\pm 0.07$ $(8.05\pm 0.71)\times 10^{-13}$ $2.24\pm 0.15$ $0.17\pm 0.10$ $(9.02\pm 0.91)\times 10^{-13}$
Table 2: List of averaged spectral parameters for three periods (PF, flare, total): photon index for power-law $\Gamma$, normalization factor for power-law $\Phi_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$), slope of the log-parabola $\alpha$, curvature log-parabola $\beta$, normalization factor for log-parabola $N_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$). The 4FGL-DR4 values are reported in the catalog (Ballet et al., 2023).
AGN gamma-ray spectra are generally well represented by either a log-parabola (LP) function or by a power-law (PL) one. The LP function is described as:
$\frac{dN}{dE}=\Phi_{0}\Big{(}\frac{E}{E_{0}}\Big{)}^{-\alpha-\beta Log(E/E_{0})}$ (1)
where the normalization factor $\Phi_{0}$ (MeV-1 cm-2 sec-1) is the flux density at $E_{0}$, $\alpha$ is the slope index, $\beta$ is the curvature, and $E_{0}$ (MeV) is the scale parameter.
The PL function, instead, is described as:
$\frac{dN}{dE}=N_{0}\,\Big{(}\frac{E}{E_{b}}\Big{)}^{-\Gamma}$ (2)
where the spectral parameters are the normalization factor $N_{0}$ in (MeV-1 cm-2 sec-1), spectral index $\Gamma$ and the energy scale $E_{b}$ (MeV).
The difference in the spectral state obtained from the analysis of the three considered periods can be appreciated by the comparison of their PL indices and LP spectral parameters with the corresponding values reported in the 4FGL-DR4 catalog  (Ballet et al., 2023), as shown in Table 2. The flare period is characterized by a spectral hardening that is detected as a decrease of power-law index and log-parabola slope, which becomes more striking at the time of peak flux. According to  Abdollahi et al. (2022), the sources are represented with a curved spectral model when $TS_{curv}=2[log\mathcal{L}_{curv}-log\mathcal{L}_{PL}]>4\,(2\sigma)$. In our case the curved spectral model is a LP, and all the three periods are characterized by a larger $TS_{curv}$ than threshold: $TS_{LP}=3.0\sigma$ for the total period, $TS_{LP}=3.2\sigma$ for the flare period, and $TS_{LP}=3.1\sigma$ for the peak flare period. Notably, the 4FGL-DR3 catalog reports a curvature significance of $TS_{LP}\sim 1.68\sigma$ for this source and the preferred spectral model is a power-law. On the contrary, the longer monitoring of 4FGL-DR4 (Ballet et al., 2023) points to a curvature significance of $TS_{LP}\sim 2.39\sigma$, supporting a curved spectrum, in agreement with our result. Likely the flaring activity increased the number of photons collected from this source allowing a more accurate reconstruction of its spectral form. For this reason, the spectrum during the flaring period is well represented by a log-parabola model, as shown in Fig. 2.
Figure 2: SED of 3C 216 during the flare period. The red line is the power-law function and the blue line is the log-parabola function whose spectral parameters are described in Table 2. The upper limit is reported when $TS\leq 10$.
Figure 3: Fermi-LAT light curves of 3C 216 for the three periods considered here with different binnings: in the top panel (blue) Peak Flare period daily binned, with dashed lightblue lines representing the days of Swift detection (2023-05-03, 04, 06, 08, 09); in the middle panel (green) flare period weekly binned, with the dotted orange line representing the first flare ($2022$-$12$-$08$) and the dashed orange line the beginning of the PF activity ($2023$-$05$-$01$), the blue area refers to the time range of the top panel; in the bottom panel (red) the total period with six months binning, the green and blue areas refer to the middle and top pannel, respectively. Upper limits are reported when $TS\leq 10$.
We also extracted the light curves, by allowing the flux normalization of the sources within $3^{\circ}$ of the target to vary freely while freezing all the other parameters from the baseline analysis, in Fig 3. The top panel displays the flux around the PF period using $1$-day time bins from $2023$-$04$-$28$ to $2023$-$05$-$10$. The middle panel shows the flare period with $1$-week time bins, and the bottom panel shows the total period with $6$-month time bins. The analysis of the flare period reveals that the flare activity achieved a first maximum on 2022-12-8, marked by the orange dotted line in the middle panel, followed by an epoch of moderate activity before the exceptional outburst occurring on 2023-05-01, highlighted by the orange dashed line. The light curve in PF the period illustrates the sharp rise of the flux leading to the flare peak (top panel of Fig. 3). The blue dashed lines indicate the Swift observations conducted in 2023.
We studied the spectral variability in relation to the flux during the flaring activity, as shown in the top panel of Fig. 4. In this short time range, the spectral curvature is not particularly significant, and we can use a PL model to obtain a more accurate evaluation of the spectral variability. This analysis aims to characterize the synchrotron emission model during the flare and its subsequent cooling phase. The plot shows the evolution of the flare’s spectral parameters from $2023$-$04$-$30$ at $18$:$00$:$00$ to $2023$-$05$-$04$ at $18$:$00$:$00$, in $12$-hour bins, following the flux increase that led to the maximum observed gamma-ray daily emission and the duration of the subsequent Swift multi-frequency follow-up. The observations are numbered to represent their chronological order. The chronological evolution is suggestive of an anticlockwise trail for the first four bins (numbered from 0 to 3), while the spectral index appears to become more stable at later times. This is consistent with the expectations implied by a single-zone SSC scenario, where the radiative efficiency and the energy distribution of the emitting particles are subject to the competing effects of a fast energy injection, occurring nearly simultaneously for all particles, and a radiative cooling that affects high energy particles more quickly than low energy ones  (Kirk et al., 1998). In the bottom panel of Fig. 4, we associated the light curve with the corresponding spectral data points to illustrate the flux variability over time.
Figure 4: Top panel: the gamma-ray photon index of 3C 216 as a function of the gamma-ray flux. Bottom panel: the light curve of 3C 216. In both panels, the numbers indicate the chronological order of emissions, with the starting point at $2023$-$04$-$30$ at $18$:$05$:$49$ designed as number 00, and the finishing point at $2023$-$05$-$04$ at $18$:$05$:$49$ designed as number $8$. The spectral index is evaluated in a 121212\,12hr binned light curve.
3 Neil Gehrels Swift XRT and UVOT data
Following the powerful gamma-ray outburst of 3C 216, detected on 2023-05-01, (Giacchino et al., 2023), a Target of Opportunity (ToO) request was sent to the Swift Gamma-Ray Burst Mission (Gehrels et al., 2004). Swift executed five visits of the target, on 2023-05-03, 04, 06, 08 and 09, accumulating observations with both the XRT and UVOT instruments. The observations were executed with a roughly regular spacing of $1.5$ days between subsequent visits. During the observations, the XRT instrument worked in photon counting mode, while UVOT performed a sequence of exposures using the $V$, $B$, $U$, $UW1$, $UM2$ and $UW2$ photometric filters.
3.1 Analysis of Swift data
3C 216 was detected in X-ray/UV/optical frequencies with values range of the X-ray flux of $\Phi_{X}\sim(2.3-3.2)\times 10^{-12}\,\mathrm{erg/cm^{2}/s}$, UV magnitudes in the range $14.6$-$15.8$, and optical ones in the range $15.2$-$16.5$. In Table 3 we report also the fluxes in different filters for UVOT, and the flux and photon index for XRT. We report also the results for archival observations carried out on 2010-10-21.
Time 2010-11-21 2023-05-03b 2023-05-04b 2023-05-06b 2023-05-08c 2023/05/09c
UVOT UVV $(5.69\pm 0.82)^{a}$ $(1.29\pm 0.02)^{b}$ $(1.20\pm 0.02)^{b}$ $(0.92\pm 0.02)^{b}$ $(6.00\pm 0.15)^{c}$ $(4.31\pm 0.11)^{c}$
UVOT UBB $(7.69\pm 0.70)^{a}$ $(1.46\pm 0.03)^{b}$ $(1.47\pm 0.03)^{b}$ $(1.10\pm 0.03)^{b}$ $(6.84\pm 0.22)^{c}$ $(4.97\pm 0.16)^{c}$
UVOT UUU $(4.92\pm 0.41)^{a}$ $(1.48\pm 0.02)^{b}$ $(1.41\pm 0.03)^{b}$ $(1.12\pm 0.02)^{b}$ $(6.70\pm 0.22)^{c}$ $(5.27\pm 0.17)^{c}$
UVOT UM2 $(1.35\pm 0.34)^{a}$ $(1.65\pm 0.03)^{b}$ $(1.60\pm 0.03)^{b}$ $(1.33\pm 0.03)^{b}$ $(8.87\pm 0.26)^{c}$ $(6.31\pm 0.16)^{c}$
UVOT UW1 $(2.38\pm 0.40)^{a}$ $(1.75\pm 0.03)^{b}$ $(1.74\pm 0.03)^{b}$ $(1.35\pm 0.03)^{b}$ $(9.45\pm 0.28)^{c}$ $(7.18\pm 0.19)^{c}$
UVOT UW2 $(2.16\pm 0.38)^{a}$ $(1.64\pm 0.04)^{b}$ $(1.57\pm 0.05)^{b}$ $(1.43\pm 0.04)^{b}$ $(9.29\pm 0.40)^{c}$ $(7.21\pm 0.29)^{c}$
XRT $(1.54\pm 0.41)^{c}$ $(3.24\pm 0.42)^{c}$ $(2.36\pm 0.42)^{c}$ $(2.40\pm 0.49)^{c}$ $(2.26\pm 0.61)^{c}$ $(2.84\pm 0.72)^{c}$
$\Gamma_{X}$ $1.24\pm 0.57$ $1.91\pm 0.11$ $2.03\pm 0.16$ $1.58\pm 0.17$ $1.62\pm 0.23$ $1.29\pm 0.18$
Table 3: Table of fluxes observed by Swift for UVOT (in six bands) and XRT (from $0.3$ to $10$ keV). Photon index $\Gamma_{X}$ for spectral model in X-ray. The error are $1\sigma$. The fluxes are in erg/cm2/sec.a is ×$\times 10^{-14}$. b is ×$\times 10^{-11}$. c is ×$\times 10^{-12}$.
The UVOT data were processed according to a standard UVOT software analysis, using heasoft-6.32.1444https://heasarc.gsfc.nasa.gov/docs/software/heasoft/. At first, we combined the exposures of each filter with uvotimsum, so to obtain one image per visit per filter. Then we extracted the flux and magnitude in each band-pass, together with their associated errors, using the task uvotdetect. The source was clearly detected in a high state, achieving a statistical significance larger than $20\sigma$ in all the visits and with all the photometric filters. The first Swift observation recorded a $U$-band magnitude of $(14.67\pm 0.02)\,$mag, more than 3 magnitudes brighter than the archival value of 17.8817.8817.88\,17.88mag reported in literature in the same band (Ryle & Sandage, 1964), while subsequent observations traced a steadily decreasing trend. In spite of the long time elapsed since this archival photometric determination, the source has also been observed more recently in ugriz photometry by the Sloan Digital Sky Survey, resulting in even fainter magnitude values in all pass-bands (Adelman-McCarthy et al., 2008), and by Swift itself in 2010, obtaining magnitudes of the order of $19$. We can therefore conclude that the UVOT magnitudes observed during the gamma-ray outburst correspond to a high state of the source. In order to recover the intrinsic state, we corrected the observed fluxes, accounting for the effects of foreground extinction due to the Milky Way’s interstellar medium. Since 3C 216 is located far away from the Galactic Plane, it is subject to a modest reddening effect, having $A_{V}=0.052$ mag, $A_{B}=0.069$ mag and $A_{U}=0.082$ mag (Schlafly & Douglas, 2011). To correct the $UV$ pass-bands, we derived the extinction coefficients for the $UW2$, $UM2$ and $UW1$ according to Roming et al. (2009) and we compared them with the corresponding values estimated by Yi et al. (2023), who claim to have updated photometric information. Since the two methods provided consistent estimates, with the largest difference being a 2% smaller intrinsic flux in the $UM2$ band according to Yi et al. (2023), with respect to Roming et al. (2009), we eventually used the corrections derived by Yi et al. (2023).
The XRT data were processed using xrtpipeline v3.7.0555https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/xrt.html. For each observation, spectral analysis of the source was performed within a $35$ arcsec radius of target, with background emission subtracted from an annular region between an inner radius $r_{in}\sim 100$ arcsec and an outer radius $r_{out}\sim 400$ arcsec, centered on the position of 3C 216. The spectra were then binned to ensure a minimum of one count per bin and modeled using the XSPEC v12.13.1e (Arnaud et al., 1999) package, employing the Cash statistic for minimization Cash (1979). Fluxes were extracted in the soft ($[0.5-2]$ keV) and hard ($[2-10]$ keV) bands. The adopted model is a power-law modified by neutral Galactic absorption $TB_{abs}\times z_{pow}$.
Figure 5: Light curves from UVOT data across six filters (top panel), XRT data in two energy ranges (medium panel), and Fermi-LAT daily binned data (bottom panel) over the same observation period (PF).
In Fig. 5 we present the light curves across different frequency bands for the PF period: the UVOT data for the six filters are shown in the top panel, the XRT data for the soft and hard bands in the middle panel, and the LAT data in the bottom panel. The observed trend indicates a stable decreasing evolution in the light curves across all spectral bands, with the exception of the hard X-ray range. This behavior is consistent with a scenario involving an initial simultaneous injection of energy into both the low frequency synchrotron component and the $\gamma$-ray IC component, followed by a radiative cooling phase. During this phase, while the overall power of the source decreases, the frequency of the IC scattered radiation also reduces, moving the high energy spectral hump into the hard X-ray regime. This effect can be clearly appreciated by looking at the evolution of the SED of 3C 216, plotted in Fig. 6, which shows that a relative increase of the hard-X ray flux agrees with the predictions of a SSC model, where a distribution of relativistic charged particles loses energy through radiative cooling (Tramacere et al., 2011).
Figure 6: SED for the broadband emission of 3C 216, spanning from UV to gamma-ray, based on five datasets observed during the PF period. The blue line represents the SSC model of high-energy photon emission computed with JetSeT. The green dots indicate the data collected by UVOT, XRT, and Fermi-LAT, along with their respective uncertainties. Upper limits are reported when $TS\leq 10$. Light gray points in the background represent archival, non-simultaneous data from literature, specifically from the following sources: 3C, NRAO, Ohio (Dixon, 1970), FIRST at VLA, Green Bank GB6 and North surveys, combined NRAO and Parkes survey, NVSS, VLBA, Planck ERCSC, PCCS1F and PCCS2F catalogs, WMAP catalog, TWOMASS and WISE infrared catalogs, USNO and SDSS2,6,7, optical surveys, Einstein and ROSAT X-ray source catalogs, Swift XRT SWXRT1 and 1SXPS catalogs (D’Elia et al., 2013; Evans et al., 2014), Chandra ACIS source catalog (Massaro et al., 2015), and the Fermi-LAT DR3 source catalog (Abdollahi et al., 2022) data. These data were extracted from the SSDC SED Builder and NED archives.
4 Discussion and conclusions
Due to its overall steep radio spectrum and relatively compact morphology, 3C 216 is classified as a CSS source. In the youth scenario (Fanti et al., 1995), these sources owe their compactness to a young age and represent an early evolutionary stage towards a fully developed radio galaxy size. CSS sources, and their even more compact (and younger) siblings (GHz Peaked Spectrum, or GPS, sources, O’Dea et al. 1991), were predicted to be good candidates for GeV emission, as they combine on sub-galactic scales both the presence of recently injected relativistic particles and abundant photon fields from the central regions of their hosts (Stawarz et al., 2008). However, only a handful of CSS/GPS sources have been detected individually (Migliori et al., 2016; Principe et al., 2020) and even a stacking analysis has not revealed a collective signal from this population (Principe et al., 2021).
Except for a few outstanding sources detected because of their extreme proximity, a significant contribution from Doppler beamed components, such as relativistic jets, is required for CSS and GPS sources to be detected in gamma rays. This is suggested by elements such as the position in the luminosity-photon index diagram, the identification with quasar hosts, and the presence of variability. While Principe et al. (2021) already pointed out all these characteristics for 3C 216, the flare analyzed here reveals a level of flux variability significantly more extreme than anything observed before in this source.
As far as the radio properties are concerned, noteworthy features include an upturn in the integrated spectrum around a few GHz (Taylor et al., 1995), the presence of a compact central component with flat spectrum and a strong bend, and the detection of superluminal motions on parsec scales (Venturi et al., 1993; Paragi et al., 2000). These factors clearly indicate that along with the non-relativistic steep spectrum lobes, which may be seen in projection, the source has a blazar core seen under a small viewing angle.
For modeling and fitting radiative emission in 3C 216, we used the JetSeT tool (Tramacere, 2020). JetSeT is an open source C/Python framework for reproducing the radiative and acceleration processes acting in relativistic jets, and Galactic objects (beamed and unbeamed), allowing numerical models to be fit to the observed data. The included radiative processes in the tool are: synchrotron self-Compton, external Compton (EC) and EC against the CMB.
We decided to analyze the time variability of the multi-frequency SED to understand the physical process of the flare activity of 3C 216. We chose to construct our base-line model with the data set corresponding to 2023-05-03, because it is the earliest time for which we obtained simultaneous multi-frequency coverage of the source. We then tested the model on the following days, to check whether the SED evolution is consistent with an initial energy injection, followed by radiative losses and cooling. UVOT flux densities and X-ray data have been already explained in the previous section. In the gamma energy range, we considered the same Fermipy parameters of Table 1 for each day.
Since the SED evolves quite coherently in nearly all the observed pass-bands, instead of considering the possibility of external contribution to the IC component, we tested whether the status of the source could be reproduced by a SSC (Jones et al., 1974). In this case the seed photons for the IC process are the synchrotron photons produced by the same population of relativistic electrons that scatter them up to gamma-ray energies. We observed that the broad band shape of the SED could be explained by a leptonic plasma blob (the emitting region) with the characteristics reported in Table 4.
Parameter Value
Size of the spherical emitting region $R=8.3\times 10^{15}\,$cm
Particle density $n=10^{3}\,\mathrm{cm^{-3}}$
Intensity of the magnetic field in $R$ $B=1\,$G
Bulk Lorentz factor in $R$ $\Gamma=8.5$
Jet viewing angle $\theta\simeq 1/\Gamma=4^{\circ}$
Table 4: Leptonic plasma blob characteristics.
The observed spectral shape of photons can be reproduced by a charged lepton population with a flat low-energy spectral index and a log-parabola energy distribution $f(\gamma)$. The required density, $N=1000\,\mathrm{cm^{-3}}$, is typical of the immediate environment surrounding an AGN, and the blob size is consistent with a variability timescale on the order of $1$ day in the observed frame. The number of emitting particles $N$ per unit volume is given by:
[…]
where we adopted a constant value of $\gamma_{min}=2$ for the minimum Lorentz factor of the emitting particles energy distribution. The spectral law is defined by a log-parabola:
$f(\gamma)=\Big{(}\frac{\gamma}{\gamma_{0}}\Big{)}^{-(s+rLog[\gamma/\gamma_{0}])}$ (4)
where $\gamma_{0}$ is the reference energy, $r$ the curvature, $s$ the spectral index.
A comparison of the adopted model with JetSeT (Massaro et al., 2006; Tramacere et al., 2009, 2011) and the broadband observed SED is illustrated in Fig. 6. Table 5 provides a summary of the changes in the spectral parameters required to account for the evolution of the SED.
Obs. Date $\gamma_{\mathrm{max}}$ $r$ $s$ $\gamma_{0}$
2023-05-03 $5\times 10^{5}$ $1.05$ $1$ 580
2023-05-04 $4\times 10^{5}$ $1.05$ $1.05$ $575$
2023-05-06 $4\times 10^{5}$ $1.07$ $1.07$ $540$
2023-05-08 $3\times 10^{5}$ $1.15$ $1.15$ $540$
2023-05-09 $3\times 10^{5}$ $1.15$ $1.2$ $530$
Table 5: List of parameter model values from JetSeT tool that evolve over time. The reported parameters include time, the maximum energy of energy interval for the emitting electron distribution ($\gamma_{\mathrm{max}}$), the curvature of log-parabola ($r$), the spectral index of log-parabola ($s$), and the reference energy of log-parabola ($\gamma_{0}$).
Interestingly, the model that fits the data of 2023-05-03 also provides an excellent framework for fitting the SED observed in the following days, just by reducing the energy of the radiating particles $\gamma_{max}$, applying a softer spectral index $s$, and a more pronounced spectral curvature $r$. Notably, we observe that the IC component in the later days enters the hard X-ray domain, leading to the observed relative enhancement of the hard X-ray flux. This supports the idea that the flaring event can confidently be interpreted as an outburst of SSC radiation, likely associated with the acceleration of a distribution of charged particles within the jet, which subsequently cooled down through radiative losses. Unfortunately, no radio data were collected during the flaring activity presented in this paper. For completeness, Fig. 6 also shows archival, non-simultaneous data extracted from the SSDC SED Builder and NED archives, covering facilities from radio to gamma-rays, to emphasize the flare process.
The tension observed between the first gamma-ray upper limit and the model of 2023-05-08 can be attributed to limited statistics, which may lead to an inappropriate representation of the spectral index in the corresponding bin. Additionally, we need to take into account the not strictly simultaneous nature of the data, since our observations in the optical, UV, and X-ray bands are representative of less than 111\,1hr of exposure each, while the gamma-ray data represent source visibility integrated over $1$ day. For the last daily bin on May 9, despite marginal agreement between the model predictions and the observations, it becomes more challenging to achieve a satisfactory fit within the framework of our simple energy-loss single-zone SSC model. This may be due to the flaring event losing power and no longer dominating over other emission regions within the source, making the single-zone emission assumption no longer suitable to describe the status of the source at later times of the event.
Figure 7: Fermi-LAT $\sqrt{TS}$ map between $100$ MeV and $300$ GeV of the region around 4FGL J0910.0+4257. In green the 3C 216 position from Petrov et al. (2005).
The simultaneous flaring activity observed in optical, UV, X-ray and gamma-rays between 2023-05-01 and 2023-05-09 provides a confirmation of the identification of the gamma-ray source with 3C 216. The TS map of 4FGL J0910.0+4257 in the total period is shown in Fig. 7. We marked in green the position of 3C 216, as reported by Petrov et al. (2005). Our results also confirm that the SSC process can well explain the production of gamma-ray outbursts from this type of radio sources. In VLBA archive [^6] https://science.nrao.edu/facilities/vlba/data-archive, there are data to study the morphology of the jet close to the central engine of the source. The results of our investigation could serve as foundation for future radio Target of Opportunity proposals if further high-energy flaring activity occurs.
The Fermi LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis. These include the National Aeronautics and Space Administration and the Department of Energy in the United States, the Commissariat à l’Energie Atomique and the Centre National de la Recherche Scientifique / Institut National de Physique Nucléaire et de Physique des Particules in France, the Agenzia Spaziale Italiana and the Istituto Nazionale di Fisica Nucleare in Italy, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), High Energy Accelerator Research Organization (KEK) and Japan Aerospace Exploration Agency (JAXA) in Japan, and the K. A. Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden. INFN and ASI personnel performed in part under ASI-INFN Agreements No. 2021-43-HH.0 The NASA Swift gamma-ray burst explorer is a MIDEX Gamma Ray Burst mission led by NASA with participation of Italy and the UK. This research has made use of the Smithsonian/NASA’s ADS bibliographic database. This research has made use of the NASA/IPAC NED database (JPL CalTech and NASA, USA). This research has made use of the archives and services of the Space Science Data Center (SSDC), a facility of the Italian Space Agency (ASI Headquarter, Rome, Italy). This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the SSDC. Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). Facilities: Fermi Gamma-ray Space TelescopeLATSwiftXRTUVOT

### v2

@@
 In this work, we perform a careful analysis of the multifrequency data (gamma-ray, X-ray, UV, optical) collected during this observational campaign.
-We find that the spectral energy distribution of the flaring source evolves in a coherent way, suggesting a common origin for the multifrequency emission.
-These results support the interpretation of the gamma-ray emission within a single zone synchrotron self-Compton (SSC) model, with important implications for the mechanisms powering high-energy radiation in AGN jets.
+We find that the spectral energy distribution of the flaring source evolves in a coherent way, supporting a common origin for the multifrequency emission.
+These results suggest that the spectral energy distribution (SED) observed during the outburst was dominated by a single emission zone, where synchrotron self-Compton processes (SSC) played a primary role.
+Since single-zone SSC models have typically less free parameters than multi-zone alternatives, they are a powerful probe of the physical conditions of the high-energy emitting regions.
+Therefore, observing SSC radiation even in CSS sources improves our understanding of the production of high-energy radiation in AGN jets.
 gamma rays: galaxies — gamma-ray astronomy — high energy astrophysics — gamma-ray sources: individual (3C 216) — radio galaxies: individual (3C 216) — X-ray sources: individual (3C 216) — blazars — relativistic jets — spectral energy distribution
@@
 1 Introduction
-The active galactic nucleus (AGN) 3C 216 ($z=0.670$, (Smith & Spinrad, 1980)) is classified as an extragalactic CSS radio source.
-On arcsecond scales, it consists of a central component surrounded by a more extended structure with an angular size of $4.5^{\prime\prime}$, corresponding to a projected linear size (LS) of $56$ kpc (Principe et al., 2021).
-Although its radio spectrum peaks at low frequency ($\nu<0.5$ GHz), an upturn is observed at a few GHz, attributed to the presence of a central flat-spectrum core (Taylor et al., 1995).
-This core exhibits a significant misalignment with the outer structure, along with superluminal jet component motion, with a velocity of approximately $\sim 4c$ (Venturi et al., 1993; Paragi et al., 2000).
-While the compact nature of CSS sources is typically attributed to their young age within an evolutionary framework (Fanti et al., 1995), the characteristics of the central component, a pronounced optical polarization and variability (Impey et al., 1991), strongly suggest the presence of a blazar core.
-The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope, is a pair-conversion telescope designed to detect photons in the energy range from $20$ MeV to $2$ TeV.
+The active galactic nucleus (AGN) 3C 216 ($z=0.670$, Smith & Spinrad 1980) is classified as an extragalactic CSS radio source.
+On arcsecond scales, it consists of a central component surrounded by a more extended halo structure with an angular size of $4.5^{\prime\prime}$, corresponding to a projected linear size (LS) of $56$ kpc (Principe et al., 2021).
+Although its radio spectrum peaks at low frequency ($\nu<0.5$ GHz), an upturn is observed at a few GHz, attributed to the presence of a central flat-spectrum core plus a faint halo component (Taylor et al., 1995).
+The core exhibits a significant misalignment with the outer structure, along with superluminal jet component motion, with a velocity of approximately $\sim 4c$ (Venturi et al., 1993; Paragi et al., 2000).
+While the compact nature of CSS sources is typically attributed to their young age within an evolutionary framework (Fanti et al., 1995), the characteristics of the central component, a pronounced optical polarization and its variability (Impey et al., 1991) strongly suggest the presence of a blazar core.
+The Large Area Telescope (LAT, Atwood et al., 2009) onboard the Fermi Gamma-ray Space Telescope, is a pair-conversion telescope designed to detect photons in the energy range from $20$ MeV to $2$ TeV.
 The LAT observed a pronounced enhancement in the gamma-ray activity from the direction of the source on 2022-12-08 (La Mura, 2022) and on 2023-05-01 (Giacchino et al., 2023).
-Multifrequency observations of the source have been carried out in order to associate the gamma-ray flare with activity at different wavelengths.
-The observations across various bands provide insights into both the origin of the gamma-ray emission and the structure of the radio jet.
-As we will demonstrate in this work, during the flare emission, the source exhibited a hard gamma-ray spectrum, in contrast to its typically soft spectrum during quiescent periods.
-This behavior can be explained by a synchrotron self-Compton (SSC) model across all frequencies, with evidence of a cooling process in the days after the flare peak.
-In the Sec.
-2 we report the main results of the Fermi-LAT observations of the flaring activity that occurred in May 2023 and in the Subsec.
-2.1 we provide a refined analysis of the Fermi-LAT data, highlighting the main features of the source in terms of its gamma-ray spectrum and light curve.
-We then describe in Sec.
-3 the Swift ToO observations of 3C 216, with a detailed analysis presented in Subsec.
-3.1.
-Finally, the discussion and conclusion are provided in Sec.
-4.
+Multifrequency observations of the source were carried out, thanks to a Swift Target of Opportunity (ToO) request, in order to associate the $\gamma$-ray flare with activity at different wavelengths.
+The observations across various bands provide insights into both the origin of the $\gamma$-ray emission and the structure of the relativistic jet.
+During the flare, the source exhibited a hard gamma-ray spectrum, as opposed to the typically soft one observed in quiescent periods, together with coherent variability from low energies up to $\gamma$ rays.
+This behavior can be explained in terms of a single-zone emission model, where it is likely that synchrotron self-Compton radiation dominated the emission across all frequencies during the outburst, with evidence of a cooling process in the following days.
+The paper is structured as follows: in Sec.
+2 we present the detection of the exceptional gamma-ray activity observed by Fermi-LAT in May 2023, with a detailed analysis of the Fermi-LAT data and its comparison with long term observations given in Sec.
+2.1; in Sec.
+3 we describe the Swift ToO observations of 3C 216, detailing their analysis in Sec.
+3.1; in Sec.
+4 we discuss our results and, finally, in Sec.
+5 we summarize our conclusions.
 2 Fermi-LAT observations
-The CSS source 3C 216 is located at R.A.
-= 137.38957 deg, DEC.
-= 42.89624 deg [J2000,  (Petrov et al., 2005)].
-It is associated with the 4FGL J0910.0+4257 source in the third data release of 4FGL catalog (4FGL-DR3, Abdollahi et al., 2022) with coordinates R.A.
-= $137.51\pm 0.11$ deg, DEC.
-= $42.96\pm 0.11$ deg.
-Fermi-LAT detected flaring activity from the direction of this source.
-A preliminary Fermi-LAT analysis of the Flare Advocate data for 2023-05-01 revealed that the daily-averaged gamma-ray flux increased by a factor of $177$ relative to the average flux reported in the 4FGL-DR3, $\langle\Phi\rangle_{\gamma}=(7.5\pm 1.4)\times 10^{-9}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ in the $1-100$ GeV energy range.
-The photon index decreased from the 4FGL averaged value, ($2.57\pm 0.10$), to ($2.11\pm 0.09$).
-This represented both the highest daily flux and the hardest photon index recorded for this source by Fermi-LAT to date (Giacchino et al., 2023).
-Due to the exceptional magnitude of this flare, the source association has been confirmed through a dedicated follow-up analysis with Fermi-LAT, complemented by multifrequency observations with Swift, which we will describe in detail later.
+3C 216 is located at $R.A.=137.38957\,$deg, $DEC.=42.89624\,$deg (J2000, Petrov et al., 2005).
+It is associated with the $\gamma$-ray source 4FGL J0910.0+4257, with coordinates $R.A.=(137.51\pm 0.11)\,$deg, $DEC.=(42.96\pm 0.11)\,$deg in the third release of the 4FGL catalog (4FGL-DR3, Abdollahi et al., 2022).
+In the course of its regular monitoring of the sky, Fermi-LAT detected an increase of activity from the direction of this source starting from the mid of November 2022.
+On 2023-05-01, a quick-look analysis showed that the source exhibited a sudden outburst, achieving a daily averaged $\gamma$-ray flux of $\langle\Phi\rangle_{daily}=(1.32\pm 0.15)\times 10^{-6}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$, a factor 176 times higher than the average flux $\langle\Phi\rangle_{\gamma}=(7.5\pm 1.4)\times 10^{-9}\,\mathrm{ph\,cm^{-2}\,s^{-1}}$ reported in the 4FGL-DR3 catalog over the $0.1-100\,$GeV energy range.
+The photon index decreased from the catalog value of $(2.52\pm 0.10)$ to $(2.11\pm 0.09)$.
+This event represented both the highest daily flux and the hardest spectral state reported for the source to date (Giacchino et al., 2023).
+Due to the exceptional magnitude of the event, multi-frequency follow-up observations were requested by activating a Swift ToO, which detected 3C 216 in a flaring state in optical, UV, and X-rays, confirming its association with 4FGL J0910.0+4257.
+We therefore analyzed the $\gamma$-ray history of the source and its associated X-ray and optical/UV observations to characterize the origin of its flaring activity.
+[…]
 2.1 Gamma-ray analysis
-The data analysis was performed with Fermipy [^1] https://fermipy.readthedocs.io/en/latest/ (version 1.2.0.) (Wood et al., 2017) and the Fermitools [^2] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software packages (version 2.2.0).
-We analyzed three periods: all Fermi-LAT data up to one month after the event (see Table 1 for details), the flare period, defined as the epoch during which Fermi-LAT detected an increasing activity trend on a weekly time scale, and the peak flare (PF) period, covering the maximum daily activity and the subsequent multi-frequency follow-up granted by the requested Swift observations.
+In order to characterize the status of 3C 216 during the outburst and to compare it with its long-term averaged properties, we analyzed the data collected by Fermi-LAT over three distinct periods.
+The first, which we refer to as the total monitoring period, covers all the Fermi-LAT observations of the source, from the start of science operations on 2008-08-04 up to 2023-06-06, approximately one month after the main $\gamma$-ray outburst.
+The second, which we call the flare period, corresponds to the start of the weekly rising trend in the activity of the source, on 2022-11-14, up to the end of the total period.
+The third one, which we call the flare peak, extends from 2023-04-28 to 2023-05-10, covering the time range corresponding to the highest level of daily $\gamma$-ray activity and the subsequent Swift follow-up campaign.
+For all these periods we performed a light curve analysis, plotted in Fig.
+1, and a standard binned analysis.
+Table 1: Table of fermipy analysis parameters.
 Parameter name Value
@@
 Isotropic diffuse model P8R3_SOURCE_V3_v1.txt
-Table 1: Table of Fermipy analysis parameters
-The time range chosen for flare period is $2022$-$11$-$14$ to $2023$-$06$-$06$, while the PF period goes from $2023$-$04$-$28$ to $2023$-$05$-$11$.
-The energy range selected for the analysis is $[0.1-300]$ GeV, using Pass 8 events (Atwood et al., 2013) and all the available photons of the SOURCE class, excluding those arriving with zenith angles greater than $90^{\circ}$ for energies smaller than $1$ GeV otherwise greater than $105^{\circ}$.
-Moreover, we choose a region of interest (ROI) of $15^{\circ}\times 15^{\circ}$ around our target, with a pixel size of $0.1^{\circ}$ and $8$ evenly spaced logarithmic energy bins.
-We utilize the P8R3_SOURCE_V3 instrumental response functions (IRFs), along with the galactic diffuse model gll_iem_v07.fits and the isotropic diffuse model P8R3_SOURCE_V3_v1.txt.
+In our study, we used fermipy v1.2.0111https://fermipy.readthedocs.io/en/latest/ (Wood et al., 2017) and the FermiTools v2.2.0.222https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ We selected all the observations in the $[0.1-300]\,$GeV energy range, using Pass 8 events (Atwood et al., 2013) and all the available photons of the SOURCE class, excluding those arriving with zenith angles greater than $90^{\circ}$ for energies smaller than $1$ GeV, otherwise greater than $105^{\circ}$.
+Moreover, we chose a region of interest (ROI) with an aperture radius of $15^{\circ}$ around the target, with a pixel size of $0.1^{\circ}$ and $8$ evenly spaced logarithmic energy bins.
+We utilized the P8R3_SOURCE_V3 instrumental response functions (IRFs), along with the galactic diffuse model gll_iem_v07.fits and the isotropic diffuse model P8R3_SOURCE_V3_v1.txt.
 The model used for the analysis includes all sources in the 4FGL-DR3 catalog [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html located at a distance $\leq 20^{\circ}$ from 4FGL J0910.0+4257.
-A summary of the parameters is reported in Table 1.
-In the spectral analysis we left the model parameters of the isotropic and diffuse background as well as the sources within $3^{\circ}$ of our target free to vary.
-To produce the spectral energy distribution (SED) of all the three periods, we used eight logarithmically spaced bins between $100$ MeV and $10$ GeV and four selected energy bins from $10$ GeV and $300$ GeV obtaining the plot in Fig.
-1.
-The plot clearly shows that the PF (blue) and flare (green) period fluxes are much higher than the values obtained considering all the Fermi-LAT (red) observations, confirming an interesting enhanced activity.
-Figure 1: SED of 3C 216 for three periods: in blue PF (Peak Flare) period, in green flare period, in red total Fermi-LAT period. The upper limit is reported when $TS\leq 10$
-Period $\Gamma$ $\Phi_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ $\alpha$ $\beta$ $N_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$
-PF $1.97\pm 0.05$ $(5.87\pm 0.46)\times 10^{-11}$ $1.79\pm 0.23$ $0.12\pm 0.09$ $(3.83\pm 0.72)\times 10^{-11}$
-Flare $2.03\pm 0.01$ $(2.45\pm 0.09)\times 10^{-11}$ $1.99\pm 0.16$ $0.06\pm 0.03$ $(1.73\pm 0.25)\times 10^{-11}$
-Total $2.17\pm 0.04$ $(1.69\pm 0.07)\times 10^{-12}$ $1.95\pm 0.20$ $0.12\pm 0.07$ $(1.02\pm 0.47)\times 10^{-12}$
-DR4 $2.43\pm 0.07$ $(8.05\pm 0.71)\times 10^{-13}$ $2.24\pm 0.15$ $0.17\pm 0.10$ $(9.02\pm 0.91)\times 10^{-13}$
-Table 2: List of averaged spectral parameters for three periods (PF, flare, total): photon index for power-law $\Gamma$, normalization factor for power-law $\Phi_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$), slope of the log-parabola $\alpha$, curvature log-parabola $\beta$, normalization factor for log-parabola $N_{0}$ ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$).
-The 4FGL-DR4 values are reported in the catalog (Ballet et al., 2023).
-AGN gamma-ray spectra are generally well represented by either a log-parabola (LP) function or by a power-law (PL) one. The LP function is described as:
+A summary of the analysis parameters is reported in Table 1.
+Figure 2: SED of 3C 216 for three periods: in flare peak (blue), in flare (green), and total (red) periods. Upper limits are reported when $TS\leq 10$.
+For the spectral analysis we left the model parameters of the isotropic and diffuse background as well as the sources within $3^{\circ}$ of our target free to vary.
+To produce the spectral energy distribution (SED) of all the three periods, we used eight logarithmically spaced bins between $100$ MeV and $10$ GeV and four energy bins from $10$ GeV and $300$ GeV, obtaining the plot in Fig.
+2.
+The result clearly shows that the flare peak (blue) and the flare (green) period fluxes are much higher than the values obtained considering all the Fermi-LAT observations (red), confirming an interesting enhanced activity.
+For the light curve analysis shown in Fig.
+1, instead, we allowed the flux normalization of the sources within $3^{\circ}$ of the target to vary freely while freezing all the other parameters from the baseline analysis.
+The top panel displays the total period with $6$-month time bins, the middle panel shows the flare period in a weekly binning, and the botton panel presents the evolution leading to the flare peak and the subsequent follow-up, using $1$-day time bins from $2023$-$04$-$28$ to $2023$-$05$-$10$.
+For the total period we have identified statistically significant variations compared to the 4FGL-DR3 averaged flux with the Bayesian Blocks algorithm (Scargle et al., 2013) as implemented in astropy [^4] https://docs.astropy.org/en/stable/api/astropy.stats.bayesian_blocks.html.
+The analysis of the flare period reveals that the flare activity achieved a first maximum on 2022-12-8, marked by the orange dotted line in the middle panel of Fig.
+1,followed by an epoch of moderate activity before the exceptional outburst occurring on 2023-05-01, highlighted by the orange dashed line.
+The blue curve in the peak period illustrates the sharp rise of the flux leading to the flare peak (bottom panel of Fig.
+1), while the light blue dashed lines indicate the Swift observations conducted during the follow-up campaign.
+The maximum flux obtained by the analysis is $\Phi_{max}=(1.22\pm 0.13)\times 10^{-6}$ ph cm-2 s-1 with a photon index of $2.06\pm 0.12$.
+The shaded areas in the top and middle panels refer to the time range of the following ones.
+Table 2: List of averaged spectral parameters for three periods (flare peak, flare, total): photon index for power-law $\Gamma$, normalization factor for power-law $\Phi_{0}$, slope of the log-parabola $\alpha$, curvature log-parabola $\beta$, normalization factor for log-parabola $N_{0}$.
+We include, for reference, also the values reported in the 4FGL-DR4 catalog (Ballet et al., 2023)
+.
+Period $\Gamma$ $\Phi_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ $\alpha$ $\beta$ $N_{0}$ $[\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}]$ Flare Peak $1.97\pm 0.05$ $(5.87\pm 0.46)\times 10^{-11}$ $1.79\pm 0.23$ $0.12\pm 0.09$ $(3.83\pm 0.72)\times 10^{-11}$ Flare $2.03\pm 0.01$ $(2.45\pm 0.09)\times 10^{-11}$ $1.99\pm 0.16$ $0.06\pm 0.03$ $(1.73\pm 0.25)\times 10^{-11}$ Total $2.17\pm 0.04$ $(1.69\pm 0.07)\times 10^{-12}$ $1.95\pm 0.20$ $0.12\pm 0.07$ $(1.02\pm 0.47)\times 10^{-12}$ 4FGL-DR4 $2.43\pm 0.07$ $(8.05\pm 0.71)\times 10^{-13}$ $2.24\pm 0.15$ $0.17\pm 0.10$ $(9.02\pm 0.91)\times 10^{-13}$
+2.2 $\gamma$-ray spectral properties
+AGN gamma-ray spectra are generally well represented by either a log-parabola (LP) function or by a power-law (PL) one. The LP function is defined as:
 $\frac{dN}{dE}=\Phi_{0}\Big{(}\frac{E}{E_{0}}\Big{)}^{-\alpha-\beta Log(E/E_{0})}$ (1)
-where the normalization factor $\Phi_{0}$ (MeV-1 cm-2 sec-1) is the flux density at $E_{0}$, $\alpha$ is the slope index, $\beta$ is the curvature, and $E_{0}$ (MeV) is the scale parameter.
-The PL function, instead, is described as:
+where the normalization factor $\Phi_{0}$ (MeV-1 cm-2 sec-1) is the flux density at $E_{0}$, $\alpha$ is the spectral index, $\beta$ is the curvature, and $E_{0}$ (MeV) is the scale parameter.
+The PL function, instead, is defined as:
 $\frac{dN}{dE}=N_{0}\,\Big{(}\frac{E}{E_{b}}\Big{)}^{-\Gamma}$ (2)
-where the spectral parameters are the normalization factor $N_{0}$ in (MeV-1 cm-2 sec-1), spectral index $\Gamma$ and the energy scale $E_{b}$ (MeV).
-The difference in the spectral state obtained from the analysis of the three considered periods can be appreciated by the comparison of their PL indices and LP spectral parameters with the corresponding values reported in the 4FGL-DR4 catalog  (Ballet et al., 2023), as shown in Table 2.
-The flare period is characterized by a spectral hardening that is detected as a decrease of power-law index and log-parabola slope, which becomes more striking at the time of peak flux.
+where the parameters are the normalization factor $N_{0}$ in ($\mathrm{MeV^{-1}\,cm^{-2}\,s^{-1}}$), the spectral index $\Gamma$ and the energy scale $E_{b}$ (MeV).
 According to  Abdollahi et al.
-(2022), the sources are represented with a curved spectral model when $TS_{curv}=2[log\mathcal{L}_{curv}-log\mathcal{L}_{PL}]>4\,(2\sigma)$.
-In our case the curved spectral model is a LP, and all the three periods are characterized by a larger $TS_{curv}$ than threshold: $TS_{LP}=3.0\sigma$ for the total period, $TS_{LP}=3.2\sigma$ for the flare period, and $TS_{LP}=3.1\sigma$ for the peak flare period.
+(2022), the sources are represented with a curved spectral model when $TS_{curv}=2[log\mathcal{L}_{curv}-log\mathcal{L}_{PL}]>4\,(=2\sigma)$.
+In our case the likelihood for curvature is evaluated by testing LP against PL spectral models.
+All the three periods are characterized by a larger $TS_{curv}$ than threshold: $TS_{LP}=3.0\sigma$ for the total period, $TS_{LP}=3.2\sigma$ for the flare period, and $TS_{LP}=3.1\sigma$ for the flare peak period.
 Notably, the 4FGL-DR3 catalog reports a curvature significance of $TS_{LP}\sim 1.68\sigma$ for this source and the preferred spectral model is a power-law.
@@
 Likely the flaring activity increased the number of photons collected from this source allowing a more accurate reconstruction of its spectral form.
-For this reason, the spectrum during the flaring period is well represented by a log-parabola model, as shown in Fig.
-2.
+As a consequence, the spectrum during the flaring period is well represented by a log-parabola model, as shown in Fig.
+3.
 Figure 3: SED of 3C 216 during the flare period.
@@
 The upper limit is reported when $TS\leq 10$.
-Figure 3: Fermi-LAT light curves of 3C 216 for the three periods considered here with different binnings: in the top panel (blue) Peak Flare period daily binned, with dashed lightblue lines representing the days of Swift detection (2023-05-03, 04, 06, 08, 09); in the middle panel (green) flare period weekly binned, with the dotted orange line representing the first flare ($2022$-$12$-$08$) and the dashed orange line the beginning of the PF activity ($2023$-$05$-$01$), the blue area refers to the time range of the top panel; in the bottom panel (red) the total period with six months binning, the green and blue areas refer to the middle and top pannel, respectively.
-Upper limits are reported when $TS\leq 10$.
-We also extracted the light curves, by allowing the flux normalization of the sources within $3^{\circ}$ of the target to vary freely while freezing all the other parameters from the baseline analysis, in Fig 3.
-The top panel displays the flux around the PF period using $1$-day time bins from $2023$-$04$-$28$ to $2023$-$05$-$10$.
-The middle panel shows the flare period with $1$-week time bins, and the bottom panel shows the total period with $6$-month time bins.
-The analysis of the flare period reveals that the flare activity achieved a first maximum on 2022-12-8, marked by the orange dotted line in the middle panel, followed by an epoch of moderate activity before the exceptional outburst occurring on 2023-05-01, highlighted by the orange dashed line.
-The light curve in PF the period illustrates the sharp rise of the flux leading to the flare peak (top panel of Fig.
-3).
-The blue dashed lines indicate the Swift observations conducted in 2023.
-We studied the spectral variability in relation to the flux during the flaring activity, as shown in the top panel of Fig.
-4.
+The evolution of the spectral state observed for the total period, the flare period and the flare peak period can be appreciated by looking at the corresponding spectral index and curvature parameters, listed in Table 2, in comparison with the values reported in 4FGL-DR4 (Ballet et al., 2023).
+We see from these results that the flare period is characterized by a spectral hardening, with respect to the average state of the source, that is detected as a decrease of the power-law index and of the log-parabola slope, becoming more evident at the time of the peak flux.
+To further characterize the spectral variability during the flare peak, we studied the spectral variability in relation to the flux, performing a 121212\,12hr binned analysis from 2023-04-30 to 2023-05-03 (Fig.
+4).
 In this short time range, the spectral curvature is not particularly significant, and we can use a PL model to obtain a more accurate evaluation of the spectral variability.
-This analysis aims to characterize the synchrotron emission model during the flare and its subsequent cooling phase.
-The plot shows the evolution of the flare’s spectral parameters from $2023$-$04$-$30$ at $18$:$00$:$00$ to $2023$-$05$-$04$ at $18$:$00$:$00$, in $12$-hour bins, following the flux increase that led to the maximum observed gamma-ray daily emission and the duration of the subsequent Swift multi-frequency follow-up.
+This analysis aims to characterize the spectral behaviour during the main $\gamma$-ray outburst and the subsequent decreasing stage.
 The observations are numbered to represent their chronological order.
-The chronological evolution is suggestive of an anticlockwise trail for the first four bins (numbered from 0 to 3), while the spectral index appears to become more stable at later times.
+In spite of the rather large uncertainties, the evolution shown in the upper panel is suggestive of an anticlockwise trail for the first four bins (numbered from 0 to 3), while the spectral index appears to become more stable at later times.
 This is consistent with the expectations implied by a single-zone SSC scenario, where the radiative efficiency and the energy distribution of the emitting particles are subject to the competing effects of a fast energy injection, occurring nearly simultaneously for all particles, and a radiative cooling that affects high energy particles more quickly than low energy ones  (Kirk et al., 1998).
@@
 Bottom panel: the light curve of 3C 216.
-In both panels, the numbers indicate the chronological order of emissions, with the starting point at $2023$-$04$-$30$ at $18$:$05$:$49$ designed as number 00, and the finishing point at $2023$-$05$-$04$ at $18$:$05$:$49$ designed as number $8$.
-The spectral index is evaluated in a 121212\,12hr binned light curve.
+In both panels, the numbers indicate the chronological order of the observations, extracted as 121212\,12hr bins starting from $2023$-$04$-$30$ at $18$:$05$:$49$ (designed as number 00) and finishing point on $2023$-$05$-$04$ at $18$:$05$:$49$ (designed as number $8$).
 3 Neil Gehrels Swift XRT and UVOT data
-Following the powerful gamma-ray outburst of 3C 216, detected on 2023-05-01, (Giacchino et al., 2023), a Target of Opportunity (ToO) request was sent to the Swift Gamma-Ray Burst Mission (Gehrels et al., 2004).
+Following the powerful gamma-ray outburst of 3C 216, detected on 2023-05-01 (Giacchino et al., 2023), a Target of Opportunity (ToO) request was sent to the Swift Gamma-Ray Burst Mission (Gehrels et al., 2004).
 Swift executed five visits of the target, on 2023-05-03, 04, 06, 08 and 09, accumulating observations with both the XRT and UVOT instruments.
@@
 During the observations, the XRT instrument worked in photon counting mode, while UVOT performed a sequence of exposures using the $V$, $B$, $U$, $UW1$, $UM2$ and $UW2$ photometric filters.
-3.1 Analysis of Swift data
-3C 216 was detected in X-ray/UV/optical frequencies with values range of the X-ray flux of $\Phi_{X}\sim(2.3-3.2)\times 10^{-12}\,\mathrm{erg/cm^{2}/s}$, UV magnitudes in the range $14.6$-$15.8$, and optical ones in the range $15.2$-$16.5$.
-In Table 3 we report also the fluxes in different filters for UVOT, and the flux and photon index for XRT.
-We report also the results for archival observations carried out on 2010-10-21.
-Time 2010-11-21 2023-05-03b 2023-05-04b 2023-05-06b 2023-05-08c 2023/05/09c
-UVOT UVV $(5.69\pm 0.82)^{a}$ $(1.29\pm 0.02)^{b}$ $(1.20\pm 0.02)^{b}$ $(0.92\pm 0.02)^{b}$ $(6.00\pm 0.15)^{c}$ $(4.31\pm 0.11)^{c}$
-UVOT UBB $(7.69\pm 0.70)^{a}$ $(1.46\pm 0.03)^{b}$ $(1.47\pm 0.03)^{b}$ $(1.10\pm 0.03)^{b}$ $(6.84\pm 0.22)^{c}$ $(4.97\pm 0.16)^{c}$
-UVOT UUU $(4.92\pm 0.41)^{a}$ $(1.48\pm 0.02)^{b}$ $(1.41\pm 0.03)^{b}$ $(1.12\pm 0.02)^{b}$ $(6.70\pm 0.22)^{c}$ $(5.27\pm 0.17)^{c}$
-UVOT UM2 $(1.35\pm 0.34)^{a}$ $(1.65\pm 0.03)^{b}$ $(1.60\pm 0.03)^{b}$ $(1.33\pm 0.03)^{b}$ $(8.87\pm 0.26)^{c}$ $(6.31\pm 0.16)^{c}$
-UVOT UW1 $(2.38\pm 0.40)^{a}$ $(1.75\pm 0.03)^{b}$ $(1.74\pm 0.03)^{b}$ $(1.35\pm 0.03)^{b}$ $(9.45\pm 0.28)^{c}$ $(7.18\pm 0.19)^{c}$
-UVOT UW2 $(2.16\pm 0.38)^{a}$ $(1.64\pm 0.04)^{b}$ $(1.57\pm 0.05)^{b}$ $(1.43\pm 0.04)^{b}$ $(9.29\pm 0.40)^{c}$ $(7.21\pm 0.29)^{c}$
-XRT $(1.54\pm 0.41)^{c}$ $(3.24\pm 0.42)^{c}$ $(2.36\pm 0.42)^{c}$ $(2.40\pm 0.49)^{c}$ $(2.26\pm 0.61)^{c}$ $(2.84\pm 0.72)^{c}$
-$\Gamma_{X}$ $1.24\pm 0.57$ $1.91\pm 0.11$ $2.03\pm 0.16$ $1.58\pm 0.17$ $1.62\pm 0.23$ $1.29\pm 0.18$
+3C 216 was detected in all the X-ray/UV/optical bands with values of the X-ray flux of the order of $\Phi_{X}\sim(2.3-3.2)\times 10^{-12}\,\mathrm{erg/cm^{2}/s}$, UV magnitudes in the range $14.6$-$15.8$, and optical ones in the range $15.2$-$16.5$.
+The results of these observations, concerning the UV, optical and X-ray fluxes, are detailed in Table 3, together with the X-ray photon indices and a comparison with a previous set of observations carried out on 2010-10-21.
 Table 3: Table of fluxes observed by Swift for UVOT (in six bands) and XRT (from $0.3$ to $10$ keV).
@@
 The error are $1\sigma$.
-The fluxes are in erg/cm2/sec.a is ×$\times 10^{-14}$.
-b is ×$\times 10^{-11}$.
-c is ×$\times 10^{-12}$.
-The UVOT data were processed according to a standard UVOT software analysis, using heasoft-6.32.1444https://heasarc.gsfc.nasa.gov/docs/software/heasoft/.
+Fluxes are in units of $10^{-12}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$
+.
+Time 2010-11-21 2023-05-03 2023-05-04 2023-05-06 2023-05-08 2023/05/09 UVOT UVV $0.06\pm 0.01$ $12.87\pm 0.18$ $11.97\pm 0.22$ $9.18\pm 0.17$ $6.00\pm 0.15$ $4.31\pm 0.11$ UVOT UBB $0.08\pm 0.01$ $14.62\pm 0.25$ $14.67\pm 0.34$ $11.00\pm 0.26$ $6.84\pm 0.22$ $4.97\pm 0.16$ UVOT UUU $0.05\pm 0.01$ $14.78\pm 0.23$ $14.12\pm 0.32$ $11.18\pm 0.23$ $6.70\pm 0.22$ $5.27\pm 0.16$ UVOT UM2 $0.01\pm 0.01$ $16.54\pm 0.28$ $15.96\pm 0.27$ $13.29\pm 0.25$ $8.87\pm 0.26$ $6.31\pm 0.16$ UVOT UW1 $0.02\pm 0.01$ $17.46\pm 0.31$ $17.40\pm 0.31$ $13.48\pm 0.29$ $9.45\pm 0.28$ $7.18\pm 0.19$ UVOT UW2 $0.02\pm 0.01$ $16.37\pm 0.39$ $15.66\pm 0.46$ $14.34\pm 0.44$ $9.30\pm 0.40$ $7.21\pm 0.29$ XRT $1.54\pm 0.41$ $3.24\pm 0.42$ $2.36\pm 0.42$ $2.40\pm 0.49$ $2.26\pm 0.61$ $2.84\pm 0.72$ $\Gamma_{X}$ $1.24\pm 0.57$ $1.91\pm 0.11$ $2.03\pm 0.16$ $1.58\pm 0.17$ $1.62\pm 0.23$ $1.29\pm 0.18$
+3.1 Analysis of Swift data
+The UVOT data were processed according to a standard UVOT software analysis, using heasoft-6.32.1555https://heasarc.gsfc.nasa.gov/docs/software/heasoft/.
 At first, we combined the exposures of each filter with uvotimsum, so to obtain one image per visit per filter.
 Then we extracted the flux and magnitude in each band-pass, together with their associated errors, using the task uvotdetect.
-The source was clearly detected in a high state, achieving a statistical significance larger than $20\sigma$ in all the visits and with all the photometric filters.
+3C 216 was clearly detected in a high state, achieving a statistical significance larger than $20\sigma$ in all the visits and with all the photometric filters.
 The first Swift observation recorded a $U$-band magnitude of $(14.67\pm 0.02)\,$mag, more than 3 magnitudes brighter than the archival value of 17.8817.8817.88\,17.88mag reported in literature in the same band (Ryle & Sandage, 1964), while subsequent observations traced a steadily decreasing trend.
@@
 Since 3C 216 is located far away from the Galactic Plane, it is subject to a modest reddening effect, having $A_{V}=0.052$ mag, $A_{B}=0.069$ mag and $A_{U}=0.082$ mag (Schlafly & Douglas, 2011).
-To correct the $UV$ pass-bands, we derived the extinction coefficients for the $UW2$, $UM2$ and $UW1$ according to Roming et al.
-(2009) and we compared them with the corresponding values estimated by Yi et al.
-(2023), who claim to have updated photometric information.
-Since the two methods provided consistent estimates, with the largest difference being a 2% smaller intrinsic flux in the $UM2$ band according to Yi et al.
-(2023), with respect to Roming et al.
-(2009), we eventually used the corrections derived by Yi et al.
+To correct the $UV$ pass-bands, we derived the extinction coefficients for the $UW2$, $UM2$ and $UW1$ according to Yi et al.
 (2023).
-The XRT data were processed using xrtpipeline v3.7.0555https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/xrt.html.
-For each observation, spectral analysis of the source was performed within a $35$ arcsec radius of target, with background emission subtracted from an annular region between an inner radius $r_{in}\sim 100$ arcsec and an outer radius $r_{out}\sim 400$ arcsec, centered on the position of 3C 216.
+The XRT data were processed using xrtpipeline v3.7.0666https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/xrt.html.
+For each observation, spectral analysis of the source was performed within a $35$ arcsec radius from the target, with background emission subtracted from an annular region between an inner radius $r_{in}\sim 100$ arcsec and an outer radius $r_{out}\sim 400$ arcsec, centered on the position of 3C 216.
 The spectra were then binned to ensure a minimum of one count per bin and modeled using the XSPEC v12.13.1e (Arnaud et al., 1999) package, employing the Cash statistic for minimization Cash (1979).
 Fluxes were extracted in the soft ($[0.5-2]$ keV) and hard ($[2-10]$ keV) bands.
-The adopted model is a power-law modified by neutral Galactic absorption $TB_{abs}\times z_{pow}$.
-Figure 5: Light curves from UVOT data across six filters (top panel), XRT data in two energy ranges (medium panel), and Fermi-LAT daily binned data (bottom panel) over the same observation period (PF).
+The adopted model is a power-law modified by neutral Galactic absorption ($TB_{abs}\times z_{pow}$).
+Figure 5: Light curves from UVOT data across six filters (top panel), XRT data in two energy ranges (medium panel), and Fermi-LAT daily binned data (bottom panel) collected during the Swift follow-up campaign
 In Fig.
-5 we present the light curves across different frequency bands for the PF period: the UVOT data for the six filters are shown in the top panel, the XRT data for the soft and hard bands in the middle panel, and the LAT data in the bottom panel.
-The observed trend indicates a stable decreasing evolution in the light curves across all spectral bands, with the exception of the hard X-ray range.
-This behavior is consistent with a scenario involving an initial simultaneous injection of energy into both the low frequency synchrotron component and the $\gamma$-ray IC component, followed by a radiative cooling phase.
-During this phase, while the overall power of the source decreases, the frequency of the IC scattered radiation also reduces, moving the high energy spectral hump into the hard X-ray regime.
-This effect can be clearly appreciated by looking at the evolution of the SED of 3C 216, plotted in Fig.
-6, which shows that a relative increase of the hard-X ray flux agrees with the predictions of a SSC model, where a distribution of relativistic charged particles loses energy through radiative cooling (Tramacere et al., 2011).
-Figure 6: SED for the broadband emission of 3C 216, spanning from UV to gamma-ray, based on five datasets observed during the PF period.
-The blue line represents the SSC model of high-energy photon emission computed with JetSeT.
+5 we present the light curves across different frequency bands for the period covered by the Swift observations: the UVOT data for the six filters are shown in the top panel, the XRT data for the soft and hard bands in the middle panel, and the LAT data in the bottom panel.
+The observed trend indicates a stable decreasing evolution in the light curves across all spectral bands, with the exception of the hard X-ray domain, which appears to be roughly constant or to exhibit a slight increase in activity.
+This behavior is consistent with the expectations of coherent emission from a single zone model.
+In this case, indeed, we would expect the energetic particles present in the jet to produce initially strong synchrotron radiation up to the UV and soft X-ray domain, while effectively scattering low energy photons up to the $\gamma$-ray domain through inverse Compton processes.
+However, as time passes and the radiating particles lose energy, the synchrotron radiation would gradually fade off and the inverse Compton scattering would rather produce hard X-ray radiation, instead of $\gamma$-rays.
+Figure 6: SED for the broadband emission of 3C 216, spanning from UV to gamma-ray, based on five datasets observed during the flare peak period.
+The blue line represents the SSC model emission computed with JetSeT.
 The green dots indicate the data collected by UVOT, XRT, and Fermi-LAT, along with their respective uncertainties.
@@
 These data were extracted from the SSDC SED Builder and NED archives.
-4 Discussion and conclusions
-Due to its overall steep radio spectrum and relatively compact morphology, 3C 216 is classified as a CSS source.
-In the youth scenario (Fanti et al., 1995), these sources owe their compactness to a young age and represent an early evolutionary stage towards a fully developed radio galaxy size.
-CSS sources, and their even more compact (and younger) siblings (GHz Peaked Spectrum, or GPS, sources, O’Dea et al.
-1991), were predicted to be good candidates for GeV emission, as they combine on sub-galactic scales both the presence of recently injected relativistic particles and abundant photon fields from the central regions of their hosts (Stawarz et al., 2008).
-However, only a handful of CSS/GPS sources have been detected individually (Migliori et al., 2016; Principe et al., 2020) and even a stacking analysis has not revealed a collective signal from this population (Principe et al., 2021).
-Except for a few outstanding sources detected because of their extreme proximity, a significant contribution from Doppler beamed components, such as relativistic jets, is required for CSS and GPS sources to be detected in gamma rays.
-This is suggested by elements such as the position in the luminosity-photon index diagram, the identification with quasar hosts, and the presence of variability.
-While Principe et al.
-(2021) already pointed out all these characteristics for 3C 216, the flare analyzed here reveals a level of flux variability significantly more extreme than anything observed before in this source.
-As far as the radio properties are concerned, noteworthy features include an upturn in the integrated spectrum around a few GHz (Taylor et al., 1995), the presence of a compact central component with flat spectrum and a strong bend, and the detection of superluminal motions on parsec scales (Venturi et al., 1993; Paragi et al., 2000).
-These factors clearly indicate that along with the non-relativistic steep spectrum lobes, which may be seen in projection, the source has a blazar core seen under a small viewing angle.
-For modeling and fitting radiative emission in 3C 216, we used the JetSeT tool (Tramacere, 2020).
-JetSeT is an open source C/Python framework for reproducing the radiative and acceleration processes acting in relativistic jets, and Galactic objects (beamed and unbeamed), allowing numerical models to be fit to the observed data.
-The included radiative processes in the tool are: synchrotron self-Compton, external Compton (EC) and EC against the CMB.
+4 Discussion
+The regular $\gamma$-ray monitoring of 3C 216, shown by Fig.
+1 suggests that the source entered a high activity state extending over several months.
+However, the outburst that resulted in the highest observed daily flux developed in a time scale of approximately 2 days and it is therefore likely to originate from a region of limited size.
+The presence of multi-frequency follow-up observations allows us to obtain a series of snap-shots of the status of the SED in the days immediately after the $\gamma$-ray record state observed on 2023-05-01 and, subsequently, to attempt an interpretation of its origin.
+To model the radiative emission, we used JetSet (Tramacere, 2020), an open source C/Python tool for reproducing the radiative and acceleration processes acting in relativistic jets and Galactic objects (beamed and unbeamed), allowing numerical models to be fit to the observed data.
+JetSet provides models to calculate SEDs obtained by combining synchrotron-self Compton and external Compton (EC), using different types of source seed radiation field, as well as the Cosmic Microwave Background (CMB).
 We decided to analyze the time variability of the multi-frequency SED to understand the physical process of the flare activity of 3C 216.
 We chose to construct our base-line model with the data set corresponding to 2023-05-03, because it is the earliest time for which we obtained simultaneous multi-frequency coverage of the source.
-We then tested the model on the following days, to check whether the SED evolution is consistent with an initial energy injection, followed by radiative losses and cooling.
-UVOT flux densities and X-ray data have been already explained in the previous section.
-In the gamma energy range, we considered the same Fermipy parameters of Table 1 for each day.
-Since the SED evolves quite coherently in nearly all the observed pass-bands, instead of considering the possibility of external contribution to the IC component, we tested whether the status of the source could be reproduced by a SSC (Jones et al., 1974).
-In this case the seed photons for the IC process are the synchrotron photons produced by the same population of relativistic electrons that scatter them up to gamma-ray energies.
-We observed that the broad band shape of the SED could be explained by a leptonic plasma blob (the emitting region) with the characteristics reported in Table 4.
+In principle, there are many possibilities that could explain the observed $\gamma$-ray emission, including hadronic and leptonic processes, contributions from external radiation fields and multi-zone emission scenarios.
+However, such models are characterized by a rather large number of free parameters and they require rather complicated fine tuning to predict a coherent SED evolution, particularly in the case of multi-zone scenarios.
+If, on the other hand, a single-zone model is to be preferred, it is very likely that the $\gamma$-ray production mechanism is IC scattering by charged relativistic particles, which will also produce strong synchrotron emission.
+We therefore decided to reproduce the first available multi-frequency SED in the framework of a SSC model (Jones et al., 1974), testing whether the SEDs observed at later times could be interpreted by the same model, after accounting for energy loss, as it would be expected in this case.
+Another advantage of the SSC scenario over other possible alternatives is the limited number of model parameters, which can be reduced to $10$ in the case of a log-parabola particle energy distribution.
+Table 4: Geometric and magnetic properties of the emitting region (fixed model parameters).
 Parameter Value
@@
 Jet viewing angle $\theta\simeq 1/\Gamma=4^{\circ}$
-Table 4: Leptonic plasma blob characteristics.
-The observed spectral shape of photons can be reproduced by a charged lepton population with a flat low-energy spectral index and a log-parabola energy distribution $f(\gamma)$.
+Our strategy consisted in dividing the model parameter in a set of $5$ fixed values, listed in Table 4 and used to describe the size of the radiating zone, its magnetic field and its bulk motion, and another set of variable values, listed in Table 5, which account for the energy evolution of the radiating particles as a function of time.
+We found that the observed spectral shape of photons can be reproduced by a charged lepton population with a flat low-energy spectral index and a log-parabola energy distribution $f(\gamma)$.
 The required density, $N=1000\,\mathrm{cm^{-3}}$, is typical of the immediate environment surrounding an AGN, and the blob size is consistent with a variability timescale on the order of $1$ day in the observed frame.
-The number of emitting particles $N$ per unit volume is given by:
+The number of emitting particles per unit volume $N$ is given by:
 […]
-where we adopted a constant value of $\gamma_{min}=2$ for the minimum Lorentz factor of the emitting particles energy distribution. The spectral law is defined by a log-parabola:
+where we adopted a constant value of $\gamma_{min}=2$ for the minimum Lorentz factor of the emitting particles energy distribution.
+The spectral distribution of particles is described by a log-parabola:
 $f(\gamma)=\Big{(}\frac{\gamma}{\gamma_{0}}\Big{)}^{-(s+rLog[\gamma/\gamma_{0}])}$ (4)
 where $\gamma_{0}$ is the reference energy, $r$ the curvature, $s$ the spectral index.
+Table 5: List of JetSet model parameter values that evolve over time.
+The reported parameters include the time, the maximum particle Lorentz factor for the energy distribution of the emitting particles ($\gamma_{\mathrm{max}}$), the curvature of log-parabola ($r$), the spectral index of log-parabola ($s$), and the reference energy of log-parabola ($\gamma_{0}$).
+The minimum particle Lorentz factor $\gamma_{\mathrm{min}}$ is assumed to be always $2.0$.
+Obs. Date $\gamma_{\mathrm{min}}$ $\gamma_{\mathrm{max}}$ $r$ $s$ $\gamma_{0}$
+2023-05-03 2.0 $5\times 10^{5}$ $1.05$ $1$ 580
+2023-05-04 2.0 $4\times 10^{5}$ $1.05$ $1.05$ $575$
+2023-05-06 2.0 $4\times 10^{5}$ $1.07$ $1.07$ $540$
+2023-05-08 2.0 $3\times 10^{5}$ $1.15$ $1.15$ $540$
+2023-05-09 2.0 $3\times 10^{5}$ $1.15$ $1.2$ $530$
 A comparison of the adopted model with JetSeT (Massaro et al., 2006; Tramacere et al., 2009, 2011) and the broadband observed SED is illustrated in Fig.
 6.
-Table 5 provides a summary of the changes in the spectral parameters required to account for the evolution of the SED.
-Obs. Date $\gamma_{\mathrm{max}}$ $r$ $s$ $\gamma_{0}$
-2023-05-03 $5\times 10^{5}$ $1.05$ $1$ 580
-2023-05-04 $4\times 10^{5}$ $1.05$ $1.05$ $575$
-2023-05-06 $4\times 10^{5}$ $1.07$ $1.07$ $540$
-2023-05-08 $3\times 10^{5}$ $1.15$ $1.15$ $540$
-2023-05-09 $3\times 10^{5}$ $1.15$ $1.2$ $530$
-Table 5: List of parameter model values from JetSeT tool that evolve over time.
-The reported parameters include time, the maximum energy of energy interval for the emitting electron distribution ($\gamma_{\mathrm{max}}$), the curvature of log-parabola ($r$), the spectral index of log-parabola ($s$), and the reference energy of log-parabola ($\gamma_{0}$).
-Interestingly, the model that fits the data of 2023-05-03 also provides an excellent framework for fitting the SED observed in the following days, just by reducing the energy of the radiating particles $\gamma_{max}$, applying a softer spectral index $s$, and a more pronounced spectral curvature $r$.
-Notably, we observe that the IC component in the later days enters the hard X-ray domain, leading to the observed relative enhancement of the hard X-ray flux.
-This supports the idea that the flaring event can confidently be interpreted as an outburst of SSC radiation, likely associated with the acceleration of a distribution of charged particles within the jet, which subsequently cooled down through radiative losses.
-Unfortunately, no radio data were collected during the flaring activity presented in this paper.
+Interestingly, the model that fits the data of 2023-05-03 also provides an excellent baseline for fitting the SED observed in the following days, except for some tension on the $\gamma$-ray emission seen on 2023-05-08, just by reducing the energy of the radiating particles $\gamma_{max}$, applying a softer spectral index $s$, and a more pronounced spectral curvature $r$.
+Notably, we observe that the IC component in the later days enters the hard X-ray domain, in agreement with the non decreasing trend seen only for this band in Fig.
+5.
+This supports the idea that the peak of the flare can be interpreted as a single-zone outburst, where SSC radiation played a dominant role, likely due to the acceleration of a distribution of charged particles within the jet, which subsequently cooled down through radiative losses.
+Unfortunately, no radio data were collected during the flaring activity presented in this paper, therefore we can neither verify nor exclude that other processes were involved, at least on the long term.
 For completeness, Fig.
-6 also shows archival, non-simultaneous data extracted from the SSDC SED Builder and NED archives, covering facilities from radio to gamma-rays, to emphasize the flare process.
-The tension observed between the first gamma-ray upper limit and the model of 2023-05-08 can be attributed to limited statistics, which may lead to an inappropriate representation of the spectral index in the corresponding bin.
-Additionally, we need to take into account the not strictly simultaneous nature of the data, since our observations in the optical, UV, and X-ray bands are representative of less than 111\,1hr of exposure each, while the gamma-ray data represent source visibility integrated over $1$ day.
+6 also shows archival, non-simultaneous data extracted from the SSDC SED Builder and NED archives, covering facilities from radio to gamma-rays, to emphasize the magnitude of the peak activity with respect to the historical properties of 3C 216.
+The tension observed between the first gamma-ray upper limit and the model of 2023-05-08 can be attributed to limited statistics, which may lead to an inappropriate representation of the spectral index in the estimate of the lowest energy $\gamma$-ray UL.
+Additionally, we need to take into account the non strictly simultaneous nature of the data, since our observations in the optical, UV, and X-ray bands are representative of less than 111\,1hr of exposure each, while the gamma-ray data represent the source visibility integrated over a full day.
 For the last daily bin on May 9, despite marginal agreement between the model predictions and the observations, it becomes more challenging to achieve a satisfactory fit within the framework of our simple energy-loss single-zone SSC model.
-This may be due to the flaring event losing power and no longer dominating over other emission regions within the source, making the single-zone emission assumption no longer suitable to describe the status of the source at later times of the event.
+This may be due to the flaring event losing power and being no longer dominant over other emission regions within the source, thus leading to a break-down of the single-zone emission assumption at later times of the event.
 Figure 7: Fermi-LAT $\sqrt{TS}$ map between $100$ MeV and $300$ GeV of the region around 4FGL J0910.0+4257. In green the 3C 216 position from Petrov et al. (2005).
-The simultaneous flaring activity observed in optical, UV, X-ray and gamma-rays between 2023-05-01 and 2023-05-09 provides a confirmation of the identification of the gamma-ray source with 3C 216.
+5 Conclusions
+The simultaneous flaring activity observed in optical, UV, X-ray and gamma-rays between 2023-05-01 and 2023-05-09 provides a confirmation of the identification of the $\gamma$-ray source 4FGL J0910.0+4257 with 3C 216.
 The TS map of 4FGL J0910.0+4257 in the total period is shown in Fig.
@@
 (2005).
-Our results also confirm that the SSC process can well explain the production of gamma-ray outbursts from this type of radio sources.
+Our results also suggest that the SSC process can well explain the production of $\gamma$-ray outbursts from this type of radio sources, thus being more universally applicable across different AGN classes than previously thought.
 In VLBA archive [^7] https://science.nrao.edu/facilities/vlba/data-archive, there are data to study the morphology of the jet close to the central engine of the source.
 The results of our investigation could serve as foundation for future radio Target of Opportunity proposals if further high-energy flaring activity occurs.
+Due to its overall steep radio spectrum and relatively compact morphology, 3C 216 is classified as a CSS source.
+In the youth scenario proposed by Fanti et al.
+(1995), these sources owe their compactness to a young age and represent an early evolutionary stage towards a fully developed radio galaxy size.
+CSS sources, and their even more compact (and younger) siblings (GHz Peaked Spectrum, or GPS, sources, O’Dea et al.
+1991), were predicted to be good candidates for GeV emission, as they combine on sub-galactic scales both the presence of recently injected relativistic particles and abundant photon fields from the central regions of their hosts (Stawarz et al., 2008).
+However, only a handful of CSS/GPS sources have been detected individually (Migliori et al., 2016; Principe et al., 2020) and even a stacking analysis has not revealed a collective signal from this population (Principe et al., 2021).
+Except for a few outstanding sources detected because of their extreme proximity, a significant contribution from Doppler beamed components, such as relativistic jets, is required for CSS and GPS sources to be detected in $\gamma$ rays.
+This is suggested by elements such as the position in the luminosity-photon index diagram, the identification with quasar hosts, and the presence of variability.
+While Principe et al.
+(2021) already pointed out all these characteristics for 3C 216, the flare analyzed here reveals a level of flux variability significantly more extreme than anything observed before in this source.
+As far as the radio properties are concerned, noteworthy features include an upturn in the integrated spectrum around a few GHz (Taylor et al., 1995), the presence of a compact central component with flat spectrum and a strong bend, and the detection of superluminal motions on parsec scales (Venturi et al., 1993; Paragi et al., 2000).
+These factors clearly indicate that along with the non-relativistic steep spectrum lobes, which may be seen in projection, the source has a blazar core seen under a small viewing angle.
+Acknowledgments
 The Fermi LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis.
@@
 Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden.
-INFN and ASI personnel performed in part under ASI-INFN Agreements No.
-2021-43-HH.0 The NASA Swift gamma-ray burst explorer is a MIDEX Gamma Ray Burst mission led by NASA with participation of Italy and the UK.
+INFN and ASI personnel performed in part under ASI-INFN Agreements No. 2021-43-HH.0.
+The NASA Swift gamma-ray burst explorer is a MIDEX Gamma Ray Burst mission led by NASA with participation of Italy and the UK.
 This research has made use of the Smithsonian/NASA’s ADS bibliographic database.
@@
 Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations).
+We also gratefully acknowledge Giacomo Principe, Filippo D’Ammando, Melissa Pesce-Rollins, Deirdre Horan, and anonymus external referees whose comments and suggestions contributed to shaping this work.
 Facilities: Fermi Gamma-ray Space TelescopeLATSwiftXRTUVOT